Constants Part A The resolution of a telescope is ultimately limited by the diameter of its objective lens or mirror. A typical amateur astronomers telescope may have a 7.0 in. -diameter mirror What is the minimum angular separation (in arc seconds) of two stars that can be resolved with a 7.0 in. scope? (Take ? to be at the center of the visible spectrum, about 550 nm, and note that 1 arc sec 1/3600 ) Express your answer using two significant figures. 0.78 arc seconds Previous Answers Correct Part B What is the minimum distance (in km) between two points on the Moons surface that can be resolved by a 7.0 in. . scope? (Note: The average distance from Earth to the Moon is 384,400 km.) Express your answer using two significant figures. km Submit
Answer #1

y = theta(in rad)*distance

theta(in rad) = 1.22 lambda/diameter

= 1.22*550*10^-9 m/(7*0.0254m) = 3.7739*10^-6 rad (you must have calculated this for part a)

Thus,

y = (3.7739*10^-6)* 384,400km

= 1.45 km = 1.5 km (2 sf)